# Understanding and Characterizing Obfuscated Funds Transfers in Ethereum Smart Contracts

### v1

@@ 5.3. Case I: MEV Bots
Due to different exchange rates across multiple decentralized exchanges (DEXs), various arbitrage opportunities exist. MEV bot contracts are thus developed and deployed to exploit the opportunities and gain profits with techniques such as front-running (Yüksel, [n. d.]), back-running (MAINAR et al., [n. d.]), and sandwich attacks (Mazorra et al., 2022). Our study finds that some MEV bot contracts leverage heavily obfuscated to hide their profit-making logic and thwart analysis. Our analysis of the highly obfuscated MEV bot contracts reveals four unique strategies.
@@
    Fallback only. Under this strategy, MEV bot contracts only implement the fallback functions to parse the calldata, which then jump to the corresponding location to continue the execution. Eliminating the 4-byte function selectors poses additional challenges for static analysis, such as function identification and call graph generation. The code typically features numerous SWAP and JUMPI pseudo-branches. This strategy is highly relevant to F3 and F4 features in Table 1 because the fallback function typically relies on external calls to handle calldata and uses conditional branches to increase the complexity.
@@
    ABI distortion. Some MEV bots manipulate function selectors by shortening or relocating them within calldata, making it difficult to identify entry points of the contract. This strategy is relevant to F1, F2, and F4 features because (1) it results in complex address generation processes, involving multiple steps; (2) manipulating ABI elements involves string operations like concatenation or hashing; and (3) it may introduce additional control flow complexity by introducing branching.
@@
    Address obfuscation. This method uses operations like PUSH4, PUSH4, and XOR to reconstruct the beneficiary address, and requires precisely-length calldata inputs, immediately reverting on mismatch. It is related to features F1 and F2 in that it introduces multiple steps to dynamically construct the transfer address and sometimes involves string manipulations. Moreover, our observation shows that it often introduces many irrelevant instructions, reducing the proportion of transfer-related instructions. Hence, it is also directly related to F6.
@@
    Runtime constraints. This strategy introduces conditional branches based on chain-specific variables (e.g., block.coinbase), preventing frontrunning in the public mempool by directing different logic flows depending on the block builder. We find that this strategy is related to F4 and F7 features since it will introduce additional conditional branches in the control flow and often emits irrelevant or misleading logs.
Figure 4. Time Series Analysis of transaction Volumes
Arbitrage transactions of MEV bots. To gain more insights into the impact of different obfuscation strategies adopted by MEV bots, we select one representative MEV bot contract for each strategy and analyze its transaction activity. We can draw several interesting observations from a time series analysis of arbitrage transactions submitted to each MEV bot, which is presented in Figure 4. Notably, the fallback‐only bot spikes to peak throughput (100 000absent100000$\approx$ 100\,000≈ 100 000 transactions) early in the period before abruptly dropping to zero, indicating a narrow exploitation window. In contrast, the ABI distortion contract ramps up gradually and sustains high volumes, while runtime constraint and address obfuscation strategies show slower growth and greater variability in transaction counts.
5.4. Case II: Ponzi Schemes
Ponzi contracts encourage users to deposit funds or purchase specific tokens by claiming high yield returns through automatic buybacks and burns, compounded mining, or cross-platform arbitrage. Then, they force participants to hold the tokens, implement multi-level commission systems, and rely on new funds from subsequent investors to support returns for earlier participants (Gazi, 2024), exhibiting classic Ponzi characteristics. When additional funds fall short, the project creator dumps tokens to reap enormous profits, triggering a collapse of the system and causing losses for participants (Ma et al., 2023).
Obfuscation strategies. Here, we use yUSDC contract, an obfuscated Ponzi contract detected by ObfProbe, to exemplify how the contract obfuscates the logic of forcing token holding, implementing buyback and burn mechanisms, and using multi-layered function wrappers.
  • Multi-level deduction and address generation. When users withdraw tokens, there is a cumbersome fee deduction process involving parameters such as devTreasury, refBonus, and buyNBurn. ObfProbe performs backward slicing from transfer operations and detects an obfuscated computational process. Furthermore, we see that the owner controls these parameter configurations and can adjust them at will.
@@ 5.4. Case II: Ponzi Schemes
Overall, we observe that Ponzi scam contracts leverage heavy obfuscation techniques to hide their core malicious logic to avoid detection.
5.5. Case III: Fake Decentralization
@@
This type of contract claims that the control of the contract is decentralized to attract participation, while maintaining obfuscated backdoor functions that allow owners to control the contract. Our study shows that two components are used to implement its malicious logic. The first one is called fake renouncement of ownership. Specifically, the contract claims that executing the renounceOwnership() can remove the centralized ownership. However, the function merely transfers the ownership to another address under the project owner’s control or simply does nothing except emit a seemingly correct log to deceive participants. The second component is malicious backdoors, which are obfuscated functions (e.g., Failsafe or Emergency). They are claimed to handle system crashes, but in fact allow the project owner to withdraw assets at any time, resulting in financial losses to participants.
@@
Obfuscation strategies. Through our detection and investigation, we see the contracts adopt three obfuscation strategies. First, the contract duplicates ownership transferring logic in multiple functions, which only differ in parameter names or variable names, to increase code complexity and hinder static analysis and manual auditing. Second, the contract often hides its core logic by inserting many empty and useless code segments, making it difficult for auditors to quickly pinpoint the core backdoor. Third, the contract publicly claims that “control has been relinquished,” while the onlyOwner modifier remains effective. Alternatively, it may contain a function (e.g., _transferOwnership(addr)), where addr is an address controlled by the owner, then the actual control is still maintained.
5.6. Case IV: Extreme Centralization
@@
Contracts in this category share a common design principle: all critical operations, from permission management to fund extraction, are ultimately controlled by a single private key, despite superficial “decentralized” interfaces or multi-role declarations. Three typical manifestations are:
@@
  • Centralized permission control. Roles such as admin, liquidateAdmin (or manager, _super, etc.) are all initialized to the deployer’s address, allowing unilateral modification of oracles, collateral ratios, fee structures, and forced liquidations.
@@
  • Arbitrarily adjustable fee/tax. They impose exorbitant buy/sell/withdrawal fees (often 10%–50%, up to 99%) via an onlyOwner-protected function, with all collected fees routed to one EOA.
@@
  • Lack of fund locking. “Staking” or “farm” contracts advertise emergency withdrawals or multi-layer strategies, but there often exist functions (e.g., emergencyWithdraw(), requestWithdraw(), or claimTokens()) that are only accessible by the owner and can enable instant drainage.
@@
Obfuscation strategies. We investigate the typical obfuscation strategies adopted by these contracts and see that they often use redundant functions, events, and misleading names to obscure their true centralization. Here, we list a few very representative ones. Please refer to Appendix A for a more comprehensive breakdown.
@@
Summary. Our study shows that it is prevalent for real-world contracts to employ various obfuscation techniques to hide their code logic, especially malicious behaviors such as MEV, Ponzi, and fake decentralization, posing huge security risks to end users.
6. Financial Impact Analysis
After understanding the prevalence and real-world usages of obfuscation techniques in smart contracts, we investigate the financial impact of obfuscation, aiming to answer RQ3. To do so, we collect a representative dataset of scam smart contracts, use ObfProbe to detect the existence of obfuscation, and qualitatively compare the financial impact between obfuscated and non-obfuscated scam smart contracts using several key metrics.
6.1. Dataset
@@
We leverage the dataset from a prior research, Li et al. (Li et al., 2023), which reports approximately 14K scam arbitrage bot contracts. To collect the ground truth information on obfuscation, we manually examine the dataset and label each contract as obfuscated or non-obfuscated. Eventually, we gain 9,197 unique obfuscated contracts and 3,826 unique contracts without obfuscation.
6.2. Comparison of Financial Loss
With the dataset, we conduct statistical analyses, visualize the inbound ETH (fund inflows) of these contracts, and calculate victim counts on the Ethereum mainnet to quantify the impact. The results indicate that obfuscated contracts exhibit a substantially more noticeable capacity for “money-grabbing” regarding economic damage and have significantly impacted users.
Comparison of Statistical Indicators. Table 6 summarizes the mean and maximum inbound ETH for both no-obfuscation and with-obfuscation groups.
Table 6. Mean and Maximum Inbound ETH
Contract Group Mean Inbound Maximum Inbound
No-Obfuscation 0.3403 ETH 83.62 ETH
With-Obfuscation 0.3454 ETH 201.74 ETH
As we can see, although the mean values are close (0.3403 ETH vs. 0.3454 ETH), the maximum inbound ETH for obfuscated contracts (201.74 ETH) is nearly 2.4× higher than that of non-obfuscated contracts (83.62 ETH). This indicates that obfuscated scams can secure significantly larger inflows, likely owing to their enhanced resistance to analysis and detection.
6.3. Time Series Analysis
To better illustrate the impact, we conduct a time series analysis on the victim counts and the aggregated inbound funds that represent the financial loss caused by the scam contracts, from 2018 to 2025, using a 15-day interval as the time unit.
Inbound ETH. As illustrated in Figure 5, the blue line representing the inbound funds of no-obfuscation scams remains near the zero axis throughout the period from 2018 to 2025, with only occasional minor increases. In contrast, the orange line, which represents obfuscation scams, displays multiple significant peaks, particularly in 2019, 2020, 2022, and 2023. These peaks highlight substantial financial losses for victims, with losses exceeding 200 ETH during the 2022-2023 period.
Figure 5. Aggregated Inbound Ether Analysis
@@ 6.3. Time Series Analysis
Victim Count. Figure 6 presents our temporal analysis of victim counts. The orange line represents the total number of victims affected by obfuscated scams, while the blue line corresponds to those impacted by non-obfuscated scams. It is evident that the number of victims from obfuscated scams consistently exceeds that of non-obfuscated scams throughout the period from 2022 to 2025. Notably, there are sharp increases during specific intervals (e.g., from the latter half of 2022 to 2023), with the total number of victims within a single 15-day period reaching several hundred. In contrast, victim counts for non-obfuscated scams remain exceptionally low during the same periods, with no significant outbreaks observed. These findings further highlight that obfuscated scams result in significantly greater harm and affect a much larger number of victims.
Figure 6. Temporal Analysis of Victim Counts
@@
Summary. Our analysis reveals that obfuscated scam contracts are more active than non-obfuscated ones and have resulted in greater financial losses and a higher number of victims. This finding aligns with our hypothesis: Obfuscation techniques enable scam contracts to operate more covertly during their initial stages, allowing them to amass larger amounts of funds and impact a greater number of victims.
7. Impact on Existing Tools
Finally, we answer RQ4 by studying how obfuscation techniques can affect the effectiveness of existing malware analysis tools. To conduct this study, we run SourceP (Lu et al., 2024), a state-of-the-art tool for detecting Ponzi schemes in smart contracts, on the same dataset described in Table 2. Then, we calculate metrics, including accuracy, recall, and F1 Score, to quantitatively compare its effectiveness in handling obfuscated and non-obfuscated samples.
Table 7. Effectiveness of SourceP for Obfuscated and Non-Obfuscated Ponzi Samples.

### v2

@@ V-C Case I: MEV Bots
Due to varying exchange rates across multiple decentralized exchanges (DEXs), arbitrage opportunities arise. MEV (Maximal Extractable Value, formerly Miner Extractable Value) refers to the value extracted by ordering, inserting, or censoring transactions within a block. MEV itself is not inherently malicious. In this work, we focus on bot behaviors whose obfuscated transfer paths and fund flows raise security concerns. MEV bot contracts are thus developed and deployed to exploit the opportunities and gain profits with techniques such as front-running [17], back-running [44], and sandwich attacks [17, 39]. Our study finds that some MEV bot contracts leverage heavily obfuscated to hide their profit-making logic and thwart analysis. Our analysis of the highly obfuscated MEV bot contracts reveals four unique MEV-specific obfuscation patterns that combine multiple obfuscation techniques from our taxonomy.
@@
    Fallback only. In this pattern, MEV bot contracts only implement the fallback functions to parse the calldata, which then jump to the corresponding location to continue the execution. Eliminating the 4-byte function selectors introduces additional challenges for static analysis, including function identification and call graph generation. The code typically features numerous SWAP and JUMPI instructions. This strategy is highly relevant to F3 and F4 features in Table I because the fallback function typically relies on external calls to handle calldata and uses conditional branches to increase the complexity. A representative MEV bot employing this obfuscation pattern can be found at address 0x6b75d8af [^00] E20b7a7ddf [^0] Ba900b4009A80.
@@
    ABI distortion. Some MEV bots manipulate function selectors by shortening or relocating them within calldata, making it difficult to identify entry points of the contract. This pattern is relevant to features F1, F2, and F4 because (1) it results in complex address generation processes, involving multiple steps; (2) manipulating ABI elements involves string operations like concatenation or hashing; and (3) it may introduce additional control flow complexity by introducing branches. A representative MEV bot employing this obfuscation pattern can be found at address 0x1F2f10d1C40777AE1da742455c65828fF36df387.
@@
    Address obfuscation. This pattern uses operations like PUSH4 and XOR to reconstruct the beneficiary address, and requires precisely-length calldata inputs, immediately reverting on mismatch. It is related to features F1 and F2 in that it introduces multiple steps to dynamically construct the transfer address and sometimes involves string manipulations. Moreover, our observation shows that it often introduces many irrelevant instructions, reducing the proportion of transfer-related instructions. Hence, it is directly captured by F1. A representative MEV bot employing this obfuscation pattern can be found at address 0xA69babEF1Ca67a37fFAf7a485DfFf3382056E78c.
@@
    Runtime constraints. This pattern introduces conditional branches based on chain-specific variables (e.g., block.coinbase), preventing frontrunning in the public mempool by directing different logic flows depending on the block builder. We find that this strategy is related to features F4 and F7, as it introduces additional conditional branches in the control flow and often emits irrelevant or misleading logs. A representative MEV bot employing this obfuscation pattern can be found at address 0x51C72848C68A965F66fA7A88855F9F7784502a7F.
Figure 5: Time Series Analysis of transaction Volumes
Arbitrage transactions of MEV bots. To gain more insights into the impact of different obfuscation patterns adopted by MEV bots, we select the representative MEV bots from each pattern to analyze their transaction activities. We draw several observations from a time-series analysis of arbitrage transactions submitted to each MEV bot, as shown in Figure 5. Notably, the fallback‐only MEV bot shows a brief burst of high activity early in the period before dropping to near-zero. The ABI distortion contract shows more sustained volume, while runtime-constraint and address-obfuscation patterns exhibit greater variability in transaction counts. These distinct profiles show that obfuscation pattern choice aligns with observable activity profiles, reinforcing the value of pattern-level analysis.
V-D Case II: Ponzi Schemes
Ponzi contracts encourage users to deposit funds or purchase specific tokens by claiming high yield returns through automatic buybacks and burns, compounded mining, or cross-platform arbitrage. Then, they force participants to hold the tokens, implement multi-level commission systems, and rely on new funds from subsequent investors to support returns for earlier participants [21], exhibiting classic Ponzi characteristics. When additional funds are short, the project creator dumps tokens to reap enormous profits, triggering a system collapse and causing losses for participants [7, 34]. We investigate obfuscated contracts and identify some representative Ponzi-specific behaviors (e.g., multi-level deduction, referral/downlines).
Obfuscation patterns. We exemplify how the contract obfuscates the logic of forcing token holding, implementing buyback and burn mechanisms, and using multi-layered function wrappers.
  • (1)
    Multi-level deduction and address generation. When users withdraw tokens, a cumbersome fee-deduction process is required, which involves parameters such as devTreasury, refBonus, and buyNBurn. ObfProbe performs backward slicing from transfer operations and detects an obfuscated computational process. Furthermore, we see that the owner controls these parameter configurations and can adjust them at will.
  • (2)
@@ V-D Case II: Ponzi Schemes
A representative contract: Deployed at 0x25cb947ebef1c56e14d5386a80262829739dbdbe on the Ethereum mainnet, the contract exhibits the above obfuscation patterns. Over 1,460 days (2020-07-29–2024-07-29), the contract involves 27,303 interactions that all transferred ETH (with 0 ERC-20 token transfer), totaling 1,433.21 ETH. This contract is found matching the following obfuscation patterns: (i) Multi-level deduction. The withdrawal path computes layered splits into devTreasury, refBonus, and buyNBurn sinks; these ratios are owner-configurable and can be adjusted at will, producing opaque fee chains. (ii) Referral/downlines. The call data encodes a referrer and users are positioned in a matrix-like structure; payouts to earlier participants are funded by subsequent deposits rather than any productive on-chain activity. (iii) Buyback/holding narrative. The code and logs contain terms such as “Burn”, “MLMReward”, and misspelled variants (e.g., “RetirementYeld”), alongside frequent “register/upgrade/reinvest/missed-earnings” events and a payout structure that redistributes funds to uplines. Through this example, we show that Ponzi contracts typically leverage heavy obfuscation techniques to hide their business logic to avoid detection. To our best knowledge, our work is the first to report this contract as a Ponzi scheme.
V-E Case III: Fake Decentralization
@@
This type of contract claims to decentralize control to attract participants, while maintaining obfuscated backdoor functions that allow owners to control the contract. Our study shows that two components are used to implement its malicious logic. The first one is called fake renouncement of ownership. Specifically, the contract states that executing renounceOwnership() can remove centralized ownership. However, the function merely transfers ownership to another address under the project owner’s control, or does nothing, except emit a seemingly correct log to deceive participants. The second component is malicious backdoors, which are obfuscated functions (e.g., Failsafe or Emergency). They are claimed to handle system crashes, but in fact allow the project owner to withdraw assets at any time, resulting in financial losses for participants.
@@
Obfuscation patterns. Through our detection and investigation, we found that this type of contract adopts three obfuscation patterns. First, the contract duplicates ownership-transfer logic across multiple functions (feature F6), which differ only in parameter or variable names, increasing code complexity and hindering static analysis and manual auditing. Second, the contract often hides its core logic by inserting many empty and useless code segments (feature F5), making it difficult for auditors to quickly pinpoint the core backdoor. Third, the contract publicly claims that “control has been relinquished,” while the onlyOwner modifier remains effective. Alternatively, it may contain a function (e.g., _transferOwnership(addr)), where addr is an address controlled by the owner, then the actual control is still maintained.
A representative contract: The above patterns are instantiated in a contract deployed at 0xb194A96AADC7e99a2462EF1669eB38E6B541DF79. It involves 11 transactions over 3 days (2020-10-06 to 2020-10-10). During this period, it receives 0 ETH and handles one ERC-20 token, YELD, with 5,0005{,}000 in and 5,0005{,}000 out (gross 10,00010{,}000, net 0). The contract is found matching the following patterns: (i) Fake renouncement: the contract claims ownership renouncement while preserving effective owner control via aliased transfer/renounce paths; (ii) Misleading logs: the contract emits “relinquished control” log events. However, the onlyOwner modifier still remains effective; (iii) Code padding/noise: the contract uses redundant stubs and near-duplicate permission paths without changing semantics.
V-F Case IV: Extreme Centralization
@@
Contracts in this category share a common design principle: all critical operations, from permission management to fund extraction, are ultimately controlled by a single address, despite superficial “decentralized” interfaces or multi-role declarations. Three typical manifestations are:
@@
  • Centralized permission control: Roles such as admin, liquidateAdmin (or manager, _super, etc.) are all initialized to the deployer’s address, allowing unilateral modification of oracles, collateral ratios, fee structures, and forced liquidations.
@@
  • Arbitrarily adjustable fee/tax: They impose exorbitant buy/sell/withdrawal fees (often 10%–50%, up to 99%) via an onlyOwner-protected function, with all collected fees routed to one address.
@@
  • Lack of funding lock. Some contracts, such as ”staking” or ”farming,” claim to have emergency withdrawals. However, there are often functions (e.g., emergencyWithdraw(), requestWithdraw(), or claimTokens()) under the owner’s control that can enable instant drainage.
@@
Obfuscation patterns. We investigate the typical obfuscation patterns adopted by these contracts and found that they often use redundant functions, events, and misleading names to obscure the centralized control, as revealed by features F6 and F7. Here, we list a few very representative ones.
  • (1)
@@
A representative contract: The above patterns are exemplified in a contract deployed at 0xd7caa679aa6e39c3891bd7a63b058bb8a269da52. Our analysis shows that it involves 439 transfer transactions within 81 days (2023-06-14 to 2023-09-03). The total received ETH is $0.2006$ ETH. The received ERC-20 tokens include 10 USDT and $19.074240298$ APE. The contract is found matching the following patterns: (i) Role masquerading: multiple roles are ultimately pointing to the same address, resulting in centralized control; (ii) Backdoored withdrawals: There are owner-controlled functions such as withdraw/withdrawErc20, which enable immediate extraction of funds in the contract.
@@
Summary. With the above four case studies, we show that it is prevalent for real-world contracts to employ various obfuscation techniques to hide their malicious behaviors such as MEV bots, Ponzi schemes, and fake decentralization, posing security risks to users.
VI Financial Impact of Smart Contract Obfuscation
Following the detailed behavioral analysis, we further investigate the financial impact of obfuscation in the real world to answer RQ3. To achieve this, we collect a representative dataset of scam smart contracts, use ObfProbe to detect obfuscation, and quantitatively compare the financial impact of obfuscated and non-obfuscated scam smart contracts to gain a deeper understanding of how obfuscation techniques affect scam financial gain. Consistent with the extreme-tail focus in Section V, we interpret impact through tail behavior rather than average losses.
VI-A Dataset
@@
We leverage the dataset from a prior study [32], which conducted a large-scale study of scams in the wild and reported approximately 13K scam contracts (which we refer to as the Li2023 dataset). To obtain ground truth on code obfuscation, we manually examine the dataset and label each contract into two groups: with-obfuscation and no-obfuscation, containing 9,197 and 3,826 unique contracts, respectively.
VI-B Overview of Financial Loss
After obtaining the two groups of contracts, we first analyze the inbound funds (fund inflows) of each group and the involved victims on the Ethereum mainnet to compare the financial loss (focus on ETH only) and then quantify the involved victims by counting unique Externally Owned Addresses (EOA) that directly send ETH to the contracts. Table XVII shows the statistics of inbound funds of each contract group. It can be seen that the average inbound funds between the two groups are very close (0.3403 ETH vs. 0.3455 ETH), and the median values are also at the same order (0.06 vs. 0.10 ETH). However, the maximum inbound funds for obfuscated contracts (201.74 ETH) is about $2.41\times$ that of non-obfuscated contracts (83.62 ETH), indicating that obfuscated contracts are more likely to cause severe financial damage to victims.
TABLE XVII: Inbound funds statistics.
Contract Group Count Inbound Funds (ETH)
Sum Median Mean Max
No-Obfuscation 3,826 1,321.82 0.10 0.3455 83.62
With-Obfuscation 9,197 3,129.997 0.06 0.3403 201.74
Although mean and median inbound funds are close, the 2.41x gap in maxima indicates that obfuscation amplifies tail losses rather than average losses, which is the more relevant signal for risk management.
VI-C Timeline Trend of Financial Loss
To better illustrate the financial impact of obfuscation, we also analyze the timeline trend of victims and the inbound funds caused by the two contract groups, from 2018 to 2025 in a 15-day interval.
Inbound Funds: Figure 6a shows the inbound funds of each contract group aggregated in the 15-day interval. It can be seen that non-obfuscated contracts (blue line) remain relatively small throughout the period from 2018 to 2025, with occasional spikes. In contrast, the inbound funds of obfuscated contracts (orange line) exhibit multiple peaks, mostly centered between 2022 and 2024. Particularly in June 2019, the 15-day aggregated inbound funds exceeded 200 ETH. Later in July 2022, the aggregated inbound funds exceeded 250 ETH. Separately, the single-contract maximum inbound funds is 201.74 ETH, as reported in Table XVII; this is a different metric from the aggregated curve. Then, in three different months from June 2022 to August 2023, the aggregated inbound funds exceeded 100 ETH each.
(a) Inbound funds aggregated by a 15-day interval.
(b) Victim EOA aggregated by a 15-day interval.
Figure 6: Time Trend Analysis
@@ VI-C Timeline Trend of Financial Loss
Victim Count. Figure 6b presents the aggregated victim addresses of each contract group in the 15-day interval. It is evident that the number of victims from obfuscated contracts (orange line) consistently exceeds that of non-obfuscated contracts (blue line) throughout the period from 2022 to 2025. Notably, there are sharp increases during specific months (e.g., from the latter half of 2022 to 2023), with the total number of victims within a single 15-day period reaching several hundred. In contrast, the number of victim addresses of non-obfuscated contracts remains much lower during the same period. Such results further highlight that obfuscated contracts can increase the likelihood of deceiving a much larger group of victims.
Scam contracts in the Li2023 dataset often rely on short-lived campaigns and strong incentives to attract participants. When transfer logic is obfuscated, the cost of manual inspection increases, amplifying the impact of these scams.
@@
Answer to RQ3. Our analysis reveals that obfuscated scam contracts are more active than non-obfuscated ones and have resulted in higher financial losses and a larger number of victims. This finding aligns with our hypothesis: Obfuscation techniques can enable scam contracts to operate more covertly during their initial stages, allowing them to generate greater profits by deceiving more victims.
VII Impact on Existing Scam Detection Tools
Finally, to answer RQ4, we examine how obfuscation affects the effectiveness of existing malware analysis tools. Note that RQ3 uses the Li2023 scam dataset, while RQ4 evaluates detection on the SourceP Ponzi-labeled set. We run SourceP [38], a representative Ponzi detector, on a single dataset: the contract set released with the SourceP paper (summarized in Table III). We additionally test MFDPonzi on the same labeled set and observe comparable precision/recall trends, so we report SourceP as the main baseline for clarity. On this dataset, we manually inspect and assign obfuscation labels according to our taxonomy (T1–T7): a contract is labeled obfuscated if at least one of T1–T7 is present; otherwise, it is labeled non-obfuscated. Ambiguous cases are excluded to avoid label noise. We then compute precision/recall/F1 to quantitatively compare SourceP’s effectiveness on obfuscated vs. non-obfuscated samples.
TABLE XVIII: Effectiveness of SourceP for obfuscated and non-obfuscated Ponzi samples.